OpenStudy (anonymous):

Find the numerical value of the expression for a)x=4 and b)x=18 \[\frac{ 5x-3 }{ 8x }\] Help please ?

OpenStudy (mathstudent55):

Let's do x = 4 together. First, rewrite the expression substituting 4 for x. Can you show what you get?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 5(4)-3 }{ 8(4) }\]

OpenStudy (mathstudent55):

Great. That is the first step. Now do the multiplications in both the numerator and denominator, and show what you get.

OpenStudy (anonymous):

i get \[\frac{ 20-3 }{ 32 }\]

OpenStudy (mathstudent55):

Correct. Again. The final step is to do the subtraction in the numerator.

OpenStudy (anonymous):

do i subract the 20-3 and divide that by 32 ?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

Just do the subtraction and write it as a fraction.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so \[\frac{ 17 }{ 32 }\]

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

That is the final answer. Now do the same procedure for x = 18

OpenStudy (anonymous):

okay got it . thanks

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

I set it up for you. If you want, just copy the pic and fill in the right side.